OpenStudy (anonymous):

27. 2 square root 5+3 square root 7 ^2 (4 points)

OpenStudy (ash2326):

Is this your question \[2\sqrt {5} + 3 \sqrt {7^2}\]?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

We have \[2\sqrt 5+ 3\sqrt {7^2}\] We know \[\sqrt {a^2}= \sqrt {a\times a}=a\] so \[2\sqrt 5+ 3 \times 7\] \[21+ 2\sqrt 5\] either we can leave here , or we can substitute the value of \(\sqrt 5 = 2.23\) so we get \[21 +2 \times 2.23=21+4.46=25.46\]